Broderick is an orthopedic​ surgeon, successfully mending broken bones and sewing together severed tendons. But this skill has allowed for Broderick to become skilled at sewing. He can sew a couture dress in​ 1/4 of the time it takes most skilled fashion designers in the industry. If Broderick earns $ 12 per minute for his orthopedic surgeon​ technique, but can earn $ 2 comma 100 for a dress he made in 4​ hours, Broderick can earn ​$ nothing more performing surgery in the same time period.
Answer #1

Earning of Broderick in making one dress = 2100$

Time taken by Broderick in making one dress = 240 minutes (4 hours)

Earning of Broderick for orthopedic surgery techniques= $12/minutes

Total earnings of Broderick for orthopedic surgery techniques in 4 hours = 12*240= 2880$

Difference in earning = 2880-2100 = $780

So Broderick can earn $780 more performing surgery in the time period of 4 hours of making one dress.